# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ Abstract
Models initialized from self-supervised pretraining may suffer from poor alignment with downstream tasks, reducing the extent to which subsequent fine-tuning can adapt pretrained features toward downstream objectives. To mitigate this, we introduce BiSSL, a novel bilevel training framework that enhances the alignment of self-supervised pretrained models with downstream tasks prior to fine-tuning. BiSSL acts as an intermediate training stage conducted after conventional self-supervised pretraining and is tasked with solving a bilevel optimization problem that incorporates the pretext and downstream training objectives in its lower- and upper-level objectives, respectively. This approach explicitly models the interdependence between the pretraining and fine-tuning stages within the conventional self-supervised learning pipeline, facilitating enhanced information sharing between them that ultimately leads to a model initialization better aligned with the downstream task. We propose a general training algorithm for BiSSL that is compatible with a broad range of pretext and downstream tasks. Using SimCLR and Bootstrap Your Own Latent to pretrain ResNet-50 backbones on the ImageNet dataset, we demonstrate that our proposed framework significantly improves accuracy on the vast majority of 12 downstream image classification datasets, as well as on object detection. Exploratory analyses alongside investigative experiments further provide compelling evidence that BiSSL enhances downstream alignment.
@@ 1 Introduction
In the absence of sufficient labeled data, self-supervised learning (SSL) has proven to be a powerful framework for training deep learning models. SSL leverages large amounts of unlabeled data to learn representations that can be effectively utilized across various downstream tasks. These representations are learned by solving a pretext task, which operates on supervisory signals extracted from the unlabeled data itself. Extensive efforts have been dedicated to designing effective pretext tasks, leading to state-of-the-art or competitive performance in various fields such as computer vision [1, 2, 3, 4, 5, 6, 7, 8], natural language processing [9, 10, 11, 12, 13] and audio signal processing [14, 15, 16, 17, 18, 19, 20]. A key advantage of SSL is that it enables the reuse of pretrained model backbones, eliminating the need to train models from scratch and thereby facilitating efficient adaptation to specific downstream tasks.
@@
Figure 1: The conventional self-supervised learning pipeline alongside the proposed pipeline involving BiSSL. The symbols ${\bf\it\theta}$ and ${\bf\it\phi}$ represent backbone and task-specific attached head parameters, respectively. When they are transmitted to the respective subsequent training stages, they are used as initializations. The objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$ represent the respective pretext pretraining and downstream fine-tuning objectives and $\mathcal{D}^{P}$, $\mathcal{D}^{D}$ the respective unlabeled pretext and labeled downstream datasets. We refer to Section 3 for further details.
When a model is pretrained on a dataset that distributionally differs from the downstream dataset, the representations learned during pretraining are likely to be misaligned with the downstream task. During fine-tuning [21, 22], this distribution misalignment is prone to causing relevant features, learned during the pretraining phase, to vanish from the representation space [23, 24, 25]. A strategy for mitigating the negative effects of these distribution discrepancies is to enhance the alignment between the self-supervised pretraining and downstream fine-tuning stages. However, since the conventional SSL pipeline treats these stages as two disjoint processes, this poses a significant challenge in devising a strategy that enhances such alignment while not compromising on the benefits that SSL offers.
@@
We suggest utilizing bilevel optimization (BLO) as a novel approach for enhancing this alignment. BLO has risen as a powerful tool for solving certain optimization problems within deep learning. It entails a main optimization problem constrained by the solution to a secondary optimization problem that depends on the parameters of the main optimization objective. This hierarchical setup causes the solutions of both optimization problems to depend on each other, either directly or implicitly, which has proven advantageous in deep learning tasks that optimize multiple inter-dependent objectives simultaneously [26]. Notable mentions of tasks within deep learning where BLO has proven useful are parameter pruning [27], invariant risk minimization [28, 29], meta-learning [30, 31], adversarial robustness [32], hyper-parameter optimization [33] and coreset selection [34].
@@
In this study, we propose BiSSL, a novel bilevel training framework that improves the alignment between the pretext pretraining and downstream fine-tuning stages in SSL. Acting as an intermediate training stage within the SSL pipeline, BiSSL frames the pretext and downstream task objectives as the lower- and upper-level objectives of a BLO problem. These objectives are connected by substituting the lower-level backbone solution into the upper-level objective. Simultaneously, the lower-level backbone is constrained to resemble the upper-level backbone parameters. This approach explicitly models the interdependence between pretext pretraining and downstream fine-tuning, enabling more effective information sharing between the pretext and downstream stages, yielding a backbone initialization better aligned with the downstream task. Figure 1 provides an overview that compares the conventional SSL pipeline with our proposed training pipeline involving BiSSL. We introduce a training algorithm that alternates between optimizing the two objectives within BiSSL, which is agnostic to the choice of pretext and downstream task. It is applied after the SSL pretraining stage, and imposes no additional assumptions about the pretraining procedure, allowing compatibility with off-the-shelf pretrained models. Through experiments using both the SimCLR [1] and Bootstrap Your Own Latent [4] pretext tasks to pretrain ResNet-50 backbones [35] on the ImageNet-1K dataset [36], we demonstrate that BiSSL significantly improves downstream performance across most of 12 downstream image classification datasets, while also significantly improving object detection accuracy. The code implementation and pretrained model weights are publicly available. [^1] https://github.com/NIPS25-16372/NIPS25_16372_BiSSL
@@ 2 Related Work
where $f$ and $g$ are referred to as the upper-level and lower-level objectives, respectively. While the lower objective $g$ has knowledge of the parameters ${\bf\it\xi}$ from the upper-level objective, the upper-level objective $f$ possesses full information of the lower objective $g$ itself through its dependence on the lower-level solution ${\bf\it\psi}^{\ast}({\bf\it\xi})$. Some works have incorporated bilevel optimization within self-supervised learning. [37] suggest formulating the contrastive self-supervised pretext task as a bilevel optimization problem, dedicating the upper-level and lower-level objectives for updating the backbone and projection head parameters respectively. Other frameworks such as the Local and Global (LoGo) [38] and Only Self-Supervised Learning (OSSL) [39] utilize auxiliary models, wherein the lower-level objective optimizes the parameters of the auxiliary model, while the upper-level objective is dedicated to training the feature extraction model. MetaMask [40] introduces a meta-learning based approach, where the upper-level learns masks that filter out irrelevant information from inputs that are provided to a lower-level self-supervised contrastive pretext task. BLO-SAM [41] is tailored towards fine-tuning the segment anything model (SAM) [42] by interchangeably alternating between learning (upper-level) prompt embeddings and fine-tuning the (lower-level) segmentation model. The aforementioned frameworks integrate bilevel optimization into either the pretraining or fine-tuning stage exclusively and are tailored towards specific pretext or downstream tasks. In contrast, our proposed BiSSL employs a BLO problem that comprehensively incorporates both training stages of pretext pretraining and downstream fine-tuning, while not being confined to any specific type of pretext or downstream task.
@@
Previous works have demonstrated that downstream performance can be enhanced by introducing techniques that modify the backbone between the pretraining and fine-tuning stages. Contrastive Initialization (COIN) [43] introduces a supervised contrastive loss to be utilized on backbones pretrained with contrastive SSL techniques. Noisy-Tune [44] perturbs the pretrained backbone with tailored noise before fine-tuning. Speaker-invariant clustering (Spin) [45] utilizes speaker disentanglement and vector quantization for improving speech representations for speech signal specific downstream tasks. RIFLE [46] conducts multiple fine-tuning sessions sequentially, where the attached downstream specific layers are re-initialized in between every session. Unlike BiSSL, these techniques either do not incorporate knowledge of both the pretext task and downstream task objectives and their relationship, or do so only implicitly.
@@ 3.1 Notation
We denote the unlabeled pretext dataset $\mathcal{D}^{P}=\{\mathbf{z}_{k}\}_{k=1}^{C_{P}}$ and labeled downstream dataset $\mathcal{D}^{D}=\{\mathbf{x}_{l},\mathbf{y}_{l}\}_{k=1}^{C_{D}}$, respectively, where $\mathbf{z}_{k},\mathbf{x}_{l}\in\mathbb{R}^{N}$. Let $f_{{\bf\it\theta}}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ denote a feature extracting backbone with trainable parameters ${\bf\it\theta}$ and $h^{T}_{{\bf\it\phi}}:\mathbb{R}^{M}\to\mathbb{R}^{Q_{T}}$ a task specific head with trainable parameters ${\bf\it{\phi}}$. Given pretext and downstream models $h^{P}_{{\bf\it\phi}_{P}}\circ f_{{\bf\it\theta}_{P}}$ and $h^{D}_{{\bf\it\phi}_{D}}\circ f_{{\bf\it\theta}_{D}}$ with ${\bf\it\theta}_{P},{\bf\it\theta}_{D}\in\mathbb{R}^{L}$, we denote the pretext and downstream training objectives $\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P};\mathcal{D}^{P})$ and $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$, respectively. To simplify notation, we omit the dataset specification from the training objectives, e.g. $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D}):=\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$.
@@ 3.2 Optimization Problem Formulation
The conventional setup of self-supervised pretraining directly followed by supervised fine-tuning relies on using a single backbone model with parameters ${\bf\it\theta}$. In that instance, we minimize $\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P})$ to produce a backbone parameter configuration ${\bf\it\theta}_{P}^{*}$ which is then used as an initialization when subsequently minimizing the downstream training objective $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$. We deviate from this by instead considering ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$ as two separate parameter vectors that are strongly correlated. In the aim of achieving an initialization ${\bf\it\theta}^{*}_{P}$ that is better aligned with the downstream task, we suggest an intermediate training stage that combines the two traditionally separate optimization problems of pretext and downstream training into a joint optimization problem through bilevel optimization, which we term BiSSL. The bilevel optimization problem is formulated as
@@
[…]
with $r$ being some convex regularization objective weighted by $\lambda\in\mathbb{R}_{+}$ enforcing similarity between ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$. The upper-level training objective in (2) is tasked with minimizing the downstream task objective $\mathcal{L}^{D}$, while the lower-level objective in (3) aims to minimize the pretext task objective $\mathcal{L}^{P}$ while also ensuring its backbone remains similar to the upper-level backbone. As seen in the left term of (2), the backbone parameters ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ are transferred into the downstream training objective, mirroring how the backbone is transferred in the conventional SSL pipeline. The second term of (2) aids convergence of the upper-level optimization during training. It also plays a crucial role in ensuring that the lower-level optimization problem (3) is non-redundant and we refer the reader to Section A.1 of Appendix A for details. Unlike the traditional SSL setup, the backbone solution of the pretext objective ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ is now a function of the parameters of the downstream backbone ${\bf\it\theta}_{D}$, since the lower-level problem is dependent on the upper-level backbone parameters.
As the upper-level objective in (2) depends on the solution ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ of the lower-level problem in (3), this enables the incorporation of information from the pretext objective when solving the upper-level optimization problem. By including a regularization objective $r$ that enforces similarity between the lower-level and upper-level backbone parameters, this setup is expected to guide the lower-level optimization toward a configuration of model backbone parameters that is more beneficial for subsequent fine-tuning on the downstream task. To more specifically understand how the pretext objective influences the downstream training procedure in this setup, we delve deeper into the gradient expression of the upper-level training objective in (2) in the following subsection.
@@ 3.3 Upper-level Derivative
Due to the dependence of the lower-level solution on the upper-level parameters, the first term of (4) includes the jacobian of the implicit function ${\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})$, referred to as the implicit jacobian (IJ). To simplify notation, we let $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$ when it is clear from context which variables are differentiated with respect to. Following an approach similar to [30], with details on the derivations and underlying assumptions outlined in Section A.2 of Appendix A, the IJ in (4) can be explicitly expressed as
$\begin{split}&\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}.\end{split}$ (5)
@@
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1},$ (6)
@@
where $I_{L}$ is the $L\crossproduct L$-dimensional identity matrix. Hence the upper-level derivative in (4) can be expressed as
@@
$\begin{split}&\frac{\text{d}F}{\text{d}{\bf\it\theta}_{D}}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1}\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})+\nabla_{{\bf\it\theta}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{D},{\bf\it\phi}_{D}).\end{split}$ (7)
@@
The inverse Hessian-vector product in (7) is computationally infeasible to calculate directly, so it is approximated using the conjugate gradient (CG) method [47, 48]. We employ a layer-wise implementation of the CG method based on that of [30] and refer to their work for more details on applying CG in a deep learning setup with BLO. While CG has proven to be an effective approach for approximating the inverse Hessian-vector products in previous works [49, 32, 30], it still introduces some computation and memory overhead due to its need for evaluations of multiple Hessian vector products. Nonetheless, the computation time of BiSSL in our experiments is merely a fraction of the self-supervised pretraining time, outlined in Section C.1 of Appendix C. Future work could explore alternative established methods for upper-level gradient approximation [26, 50, 51] that offer greater computational efficiency without compromising downstream task performance.
@@
With an explicit expression of the IJ in (6), we can interpret the impact of the scaling factor $\lambda$ from (3) and (7): When $\lambda$ is very large, the dependence of the lower-level objective on the upper-level parameters ${\bf\it\theta}_{D}$ is also very large. This effectively drives the lower-level backbone parameters toward the trivial solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})={\bf\it\theta}_{D}$. Meanwhile, the IJ in (6) approximately equals $I_{L}$, thereby diminishing the influence of the lower-level objective on the upper-level gradient in (7). This roughly makes the task of the upper-level equivalent to conventional fine-tuning. Conversely, if $\lambda$ is very small, the lower-level objective in (3) effectively defaults to conventional pretext task training. Additionally, the implicit jacobian in (6) would consist of numerically tiny entries, making the optimization of the first term in the upper-level objective in (2) equivalent to probing of the downstream head on the frozen pretext backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$.

### v5

@@ Abstract
Models initialized from self-supervised pretraining may suffer from poor alignment with downstream tasks, limiting the extent to which subsequent fine-tuning can adapt relevant representations acquired during the pretraining phase. To mitigate this, we introduce BiSSL, a novel bilevel training framework that enhances the alignment of self-supervised pretrained models with downstream tasks by explicitly incorporating both the pretext and downstream tasks into a preparatory training stage prior to fine-tuning. BiSSL solves a bilevel optimization problem in which the lower-level adheres to the self-supervised pretext task, while the upper-level encourages the lower-level backbone to align with the downstream objective. The bilevel structure facilitates enhanced information sharing between the tasks, ultimately yielding a backbone model that is more aligned with the downstream task, providing a better initialization for subsequent fine-tuning. We propose a general training algorithm for BiSSL that is compatible with a broad range of pretext and downstream tasks. We demonstrate that our proposed framework significantly improves accuracy on the vast majority of a broad selection of image-domain downstream tasks, and that these gains are consistently retained across a wide range of experimental settings. In addition, exploratory alignment analyses further underpin that BiSSL enhances downstream alignment of pretrained representations.
@@ 1 Introduction
Figure 1: BiSSL introduces an intermediate training stage conducted after pretext pretraining. The symbols $\bf\it\theta$} and $\bf\it\phi$} represent backbone and task-specific attached head parameters, respectively. When they are transmitted to the respective subsequent training stages, they are used as initializations. The objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$ represent the respective pretext pretraining and downstream fine-tuning objectives and $\mathcal{D}^{P}$, $\mathcal{D}^{D}$ the respective unlabeled pretext and labeled downstream datasets. We refer to Section 3 for further details.
Self-supervised learning (SSL) leverages large amounts of heterogeneous unlabeled data to solve pretext tasks that encourage learning of general-purpose representations, which can be subsequently repurposed to solve downstream tasks. This has lead to strong performance across domains such as vision (Chen et al., 2020; Bardes et al., 2022; He et al., 2020; Grill et al., 2020; Caron et al., 2020; 2021; He et al., 2022; Oquab et al., 2024), language (Devlin et al., 2019; Lewis et al., 2019; Brown et al., 2020; He et al., 2021; Touvron et al., 2023), and audio (Schneider et al., 2019; Baevski et al., 2020; Hsu et al., 2021; Niizumi et al., 2021; Chung and Glass, 2018; Chung et al., 2019; Yadav et al., 2024). However, downstream tasks often involve narrower domains that demand more specialized features, which can result in the self-supervised pretrained representations being misaligned with these tasks. Pretrained features that benefit the downstream task may therefore initially appear irrelevant and thus risk being degraded or overwritten during fine-tuning, thereby constraining the achievable performance (Kumar et al., 2022; Zaiem et al., 2024; Liao et al., 2025; Wang et al., 2024; Boschini et al., 2022; Hinton and Salakhutdinov, 2006). Enhancing the alignment between the pretext and downstream tasks could address this issue, allowing the pretrained representations to be more optimally utilized during fine-tuning.
@@
We suggest utilizing bilevel optimization (BLO) as a novel and effective approach for enhancing this alignment. BLO entails a main optimization problem constrained by the solution to a secondary problem that depends on the main problem’s parameters. This hierarchical setup captures and leverages interdependencies between problems and has shown strong potential in deep learning settings that involve joint optimization of multiple coupled objectives (Zhang et al., 2023a) such as parameter pruning (Zhang et al., 2022b), invariant risk minimization (Arjovsky et al., 2019; Zhang et al., 2023b), meta-learning (Rajeswaran et al., 2019; Finn et al., 2017), adversarial robustness (Zhang et al., 2021), hyper-parameter optimization (Franceschi et al., 2018), and coreset selection (Borsos et al., 2020).
@@
In this study, we propose BiSSL, a novel bilevel training framework that improves the alignment of self-supervised pretrained backbone models with downstream tasks. BiSSL introduces a preparatory training stage prior to fine-tuning that solves a bilevel optimization problem in which the pretext and downstream task objectives serve as the lower- and upper-level objectives, respectively. This bilevel setup facilitates enhanced information sharing, enabling the upper-level to guide the pretext task optimization in refining its representations to better align with the downstream objective, ultimately providing a more effective backbone initialization for fine-tuning. Figure 1 depicts how BiSSL integrates into the standard SSL pipeline. We introduce a training algorithm that alternates between optimizing the two objectives within BiSSL, which is agnostic to the choice of pretext and downstream task. Our method makes no additional assumptions about how models are pretrained, allowing compatibility with standard off-the-shelf pretrained backbones. Through experiments using both the SimCLR (Chen et al., 2020) and Bootstrap Your Own Latent (Grill et al., 2020) pretext tasks to pretrain ResNet-50 backbones (He et al., 2016) on the ImageNet-1K dataset (Deng et al., 2009), we demonstrate that BiSSL significantly improves downstream performance across most of 12 downstream image classification datasets, while also significantly improving accuracy within object detection and semantic segmentation. A suite of extended evaluations further shows that the gains remain consistent, and exploratory alignment studies of backbones highlight that BiSSL improves downstream alignment. Code and pretrained model weights are publicly available at https://github.com/GustavWZ/bissl/.
@@ 2 Related Work
where $f$ and $g$ are referred to as the upper-level and lower-level objectives, respectively. While the lower objective $g$ has knowledge of the parameters $\bf\it\xi$} from the upper-level objective, the upper-level objective $f$ possesses full information of the lower objective $g$ itself through its dependence on the lower-level solution ${\bf\it\psi}^{\ast}({\bf\it\xi})$. Some works have incorporated bilevel optimization within self-supervised learning. Gupta et al. (2022) suggest formulating the contrastive self-supervised pretext task as a bilevel optimization problem, dedicating the upper-level and lower-level objectives for updating the backbone and projection head parameters respectively. Other frameworks such as the Local and Global (LoGo) (Zhang et al., 2022a) and Only Self-Supervised Learning (OSSL) (Boonlia et al., 2022) utilize auxiliary models, wherein the lower-level objective optimizes the parameters of the auxiliary model, while the upper-level objective is dedicated to training the feature extraction model. MetaMask (Li et al., 2022b) introduces a meta-learning based approach, where the upper-level learns masks that filter out irrelevant information from inputs that are provided to a lower-level self-supervised contrastive pretext task. BLO-SAM (Zhang et al., 2024) is tailored towards fine-tuning the segment anything model (SAM) (Kirillov et al., 2023) by interchangeably alternating between learning (upper-level) prompt embeddings and fine-tuning the (lower-level) segmentation model. The aforementioned frameworks integrate bilevel optimization into either the pretraining or fine-tuning stage exclusively and are tailored towards specific pretext or downstream tasks. In contrast, our proposed BiSSL employs a BLO problem that comprehensively incorporates both training stages of pretext pretraining and downstream fine-tuning, while not being confined to any specific type of pretext or downstream task.
@@
Previous works have also demonstrated that downstream performance can be enhanced by introducing techniques that modify the backbone between the pretraining and fine-tuning stages. Contrastive Initialization (COIN) (Pan et al., 2022) introduces a supervised contrastive loss to be utilized on backbones pretrained with contrastive SSL techniques. Noisy-Tune (Wu et al., 2022) perturbs the pretrained backbone with tailored noise before fine-tuning. Speaker-invariant clustering (Spin) (Chang et al., 2023) utilizes speaker disentanglement and vector quantization for improving speech representations for speech signal specific downstream tasks. RIFLE (Li et al., 2020) conducts multiple fine-tuning sessions sequentially, where the attached downstream specific layers are re-initialized in between every session. Unlike BiSSL, these techniques either do not incorporate knowledge of both the pretext task and downstream task objectives and their relationship, or do so only implicitly.
@@ 3.1 Notation
We denote the unlabeled pretext dataset $\mathcal{D}^{P}=\{\mathbf{z}_{k}\}_{k=1}^{|\mathcal{D}^{P}|}$ and labeled downstream dataset $\mathcal{D}^{D}=\{\mathbf{x}_{l},\mathbf{y}_{l}\}_{l=1}^{|\mathcal{D}^{D}|}$, respectively, where $\mathbf{z}_{k},\mathbf{x}_{l}\in\mathbb{R}^{N}$. Let $f_{{\bf\it\theta}}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ denote a feature extracting backbone with trainable parameters $\bf\it\theta$} and $h^{T}_{{\bf\it\phi}}:\mathbb{R}^{M}\to\mathbb{R}^{Q_{T}}$ a task specific head with trainable parameters $\bf\it{\phi}$}. Given pretext and downstream models $h^{P}_{{\bf\it\phi}_{P}}\circ f_{{\bf\it\theta}_{P}}$ and $h^{D}_{{\bf\it\phi}_{D}}\circ f_{{\bf\it\theta}_{D}}$ with ${\bf\it\theta}_{P},{\bf\it\theta}_{D}\in\mathbb{R}^{L}$, we denote the pretext and downstream training objectives $\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P};\mathcal{D}^{P})$ and $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$, respectively. To simplify notation, we omit the dataset specification from the training objectives, e.g. $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D}):=\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D};\mathcal{D}^{D})$.
@@ 3.2 Optimization Problem Formulation
In standard SSL, a single backbone with parameters $\bf\it\theta$} is first trained to minimize the pretext objective $\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})$, yielding a parameter configuration $\bf\it\theta$$^{*}$ that subsequently is used as an initialization for fine-tuning on the downstream objective $\mathcal{L}^{D}({\bf\it\theta},{\bf\it\phi}_{D})$. To better align the self-supervised pretrained initialization with the downstream task, we employ BiSSL as a preparatory training stage to fine-tuning that incorporates the pretext and downstream task objectives into a joint bilevel optimization problem. BiSSL uses two distinct sets of backbone parameters ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$, corresponding to the pretext and downstream tasks, that are treated as separate but strongly correlated parameter vectors. The bilevel optimization problem is formulated as
@@
 $\begin{split}&\text{ s.t.}\,\,\boldsymbol{\theta}_{P}^{\ast}\left(\boldsymbol{\theta}_{D}\right)\in\underset{\boldsymbol{\theta}_{P}}{\text{argmin}}\,\underset{\boldsymbol{\phi}_{P}}{\text{min}}\,\mathcal{L}^{P}\left(\boldsymbol{\theta}_{P},\boldsymbol{\phi}_{P}\right)+\lambda r(\boldsymbol{\theta}_{P},\boldsymbol{\theta}_{D}),\end{split}$ (3)
with $r$ being some convex regularization objective weighted by $\lambda\in\mathbb{R}_{+}$ enforcing similarity between ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$. The upper-level training objective in (2) is tasked with minimizing the downstream task objective $\mathcal{L}^{D}$ and the lower-level objective in (3) aims to minimize the pretext task objective $\mathcal{L}^{P}$ while also ensuring its backbone remains similar to the upper-level backbone. As seen in the left term of (2), the backbone parameters ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$ are substituted into the downstream training objective, resembling how backbones conventionally are transferred after self-supervised pretraining. The lower-level’s dependence on the regularization objective $r$ enables the upper-level to observe how changes in its own backbone parameters ${\bf\it\theta}_{D}$ influence the lower-level solution ${\bf\it\theta}^{*}_{P}({\bf\it\theta}_{D})$, thereby allowing the upper-level to indirectly guide the pretext task optimization toward a lower-level backbone that better complies with the downstream task. The second term of (2) aids convergence of the upper-level objective during training by acting as a regularizer that constrains ${\bf\it\theta}_{D}$ to itself also align with the downstream task. Without this term, ${\bf\it\theta}_{D}$ would function only as an indirect controller of the lower-level, potentially leading to degenerate solutions.
@@ 3.3 Upper-level Derivative
Due to the dependence of the lower-level solution on the upper-level parameters, the first term of (4) includes the Jacobian of the implicit function ${\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})$, referred to as the implicit Jacobian (IJ). To simplify notation, we let $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$ when it is clear from context which variables are differentiated with respect to. Following an approach similar to Rajeswaran et al. (2019), with details on the derivations and underlying assumptions outlined in Section A.1, the IJ in (4) can be explicitly expressed as
$\begin{split}&\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\theta}_{D})\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\theta}_{D}))]^{-1}.\end{split}$ (5)
@@
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}_{P}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1},$ (6)
@@
where $I_{L}$ is the $L\crossproduct L$-dimensional identity matrix. Hence the upper-level derivative (4) can be expressed as
@@
$\begin{split}&\frac{\text{d}F}{\text{d}{\bf\it\theta}_{D}}=\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}_{P}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+I_{L}]^{-1}\nabla_{{\bf\it\theta}_{P}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})+\nabla_{{\bf\it\theta}_{D}}\mathcal{L}^{D}\quantity({\bf\it\theta}_{D},{\bf\it\phi}_{D}).\end{split}$ (7)
As shown in (7), the pretext objective is explicitly incorporated into the upper-level gradient, causing the upper-level updates to be influenced by how its changes affect the lower-level.
@@
The inverse Hessian-vector product in (7) is computationally infeasible to calculate directly, so it is approximated using the conjugate gradient (CG) method (Nazareth, 2009; Shewchuk, 1994). We employ a layer-wise implementation of the CG method based on that of Rajeswaran et al. (2019) and refer to their work for more details on applying CG in a deep learning setup with BLO. While CG has proven to be an effective approach for approximating the inverse Hessian-vector products in previous works (Pedregosa, 2016; Zhang et al., 2021; Rajeswaran et al., 2019), it still introduces some computation and memory overhead due to its need for evaluations of multiple Hessian vector products. Future work could explore alternative established methods for upper-level gradient approximation (Zhang et al., 2023a; Choe et al., 2023; Liu et al., 2024) that offer greater computational efficiency. We report the computation times of BiSSL alongside along pretraining and fine-tuning and discuss the Hessian vector product calculation overhead in Section C.1.
@@
With an explicit expression of the IJ in (6), we can interpret the impact of the scaling factor $\lambda$ from (3) and (7): When $\lambda$ is very large, the dependence of the lower-level objective on the upper-level parameters ${\bf\it\theta}_{D}$ is correspondingly very large. This effectively drives the lower-level backbone parameters toward the trivial solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})={\bf\it\theta}_{D}$. Meanwhile, the IJ in (6) approximately equals $I_{L}$, thereby diminishing the influence of the lower-level objective on the upper-level gradient in (7). This roughly makes the task of the upper-level equivalent to conventional fine-tuning. Conversely, if $\lambda$ is very small, the lower-level objective in (3) effectively defaults to conventional pretext task training. Additionally, the implicit Jacobian in (6) would consist of numerically tiny entries, making the optimization of the first term in the upper-level objective in (2) equivalent to linear probing of the downstream head on the frozen pretext backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$. Thus, an intermediate value of $\lambda$ is expected to yield an optimization that balances the contribution from both levels.